OpenStudy (loser66):

Given a function on the xy plane. f(x,y) such that f(2,2) = 20 and \(\dfrac{\partial f}{\partial x}=3x^2+3y\) , \(\dfrac{\partial f}{\partial y}= 3x+2y\). Then f(0,0)=?? Please, help

OpenStudy (loser66):

My attempt: \(f(0,0) -f(2,2) = \triangle f = 3x^2+3y+3x+2y\) At (2,2) , \(\triangle f = 28\), hence f(0,0) = 48, But it is very wrong. hahaha.... I don't know how to fix

ganeshie8 (ganeshie8):

use approximation formula ? looks im getting \(\Delta f \approx f_x\Delta x+f_y\Delta y = (18)*(-2) + (10)*(-2) = -56\)

OpenStudy (loser66):

The options are 1) 0 2) -1 3) -2 4) -3 5) -4

ganeshie8 (ganeshie8):

\(f(0,0) =f(2,2)+\Delta f \approx 20 - 56 = -36\)

OpenStudy (loser66):

Another thought: \(\dfrac{\partial f}{\partial x}=3x^2 +3y\\f(x,y)= x^3+3xy+h(y)\\\dfrac{\partial f}{\partial y}= 3x+h'(y) = 3x+2y\\h'(y) = 2y\\h(y) =y^2+C\) Hence \(f(x,y) = x^3+3xy +y^2+C \\f(2,2) = 2^3+3*2*2+2^2+C=20\\C= -4\) therefore \(f(x,y) = x^3+3xy+y^2-4\\f(0,0) =-4\)

OpenStudy (astrophysics):

Interesting, so that's like finding the potential function haha.

OpenStudy (loser66):

Yes, because from approximating method, I don't know how to check whether my logic is right or wrong.

OpenStudy (astrophysics):

Yeah, that method looked weird to me

OpenStudy (astrophysics):

-4 looks good but I'm not sure...@ganeshie8

OpenStudy (astrophysics):

@ganeshie8

OpenStudy (loser66):

Thank you, let's wait for him.

ganeshie8 (ganeshie8):

-36 is correct using approximation formula -4 is correct by solving the function depends on what you want

OpenStudy (loser66):

hey, no way!!! -4 is ffffffffffffffar different from -36

ganeshie8 (ganeshie8):

Thats the reason it is called "approximation" formula

OpenStudy (loser66):

but the error is so nonsense. If it is -4 and -4.5, I am ok.

ganeshie8 (ganeshie8):

You just got lucky with the present problem because there exists a potential function, this is not the case with most functions and you will end up using approximation formula

OpenStudy (loser66):

I know!! that is why I posted the question here. hehehe...

OpenStudy (loser66):

Unfortunately, when using app. and the answer is -36, how can I check the correct choice? I don't have -36 options.

ganeshie8 (ganeshie8):

how did u get the idea of finding f(x,y) directly instead of using the approximation formula ?

OpenStudy (loser66):

I don't get what you mean by "how". I try all the ways out. I saw partial derivative and I linked it to DE. hehehe...

ganeshie8 (ganeshie8):

Hmm try this : Given a function on the xy plane. f(x,y) such that f(2,2) = 20 and \(\dfrac{\partial f}{\partial x}=3x^2+3y\) , \(\dfrac{\partial f}{\partial y}= x+2y\). Then f(0,0)=??

ganeshie8 (ganeshie8):

If you think of that as vector field, notice that the curl is nonzero. So potential function cannot exist. Let me know if you ever get successful finding \(f(x,y)\) :P

OpenStudy (loser66):

as above, f(x,y) = x^3 + 3xy + h(y) f_y = 3x + h'(y) = x+2y h'(y) = -2x+2y h(y) -2xy +y^2 +C f(x,y) =x^3+3xy-2xy+y^2 +C = x^3 +xy +y^2 +C f(2,2)= 8+4+4 +C=20 C= 4 f(x,y) = x^3+xy +y^2 +4 f(0,0) =4

ganeshie8 (ganeshie8):

f(x,y) = x^3+xy +y^2 +4 f_x = 3x^2 + y this is wrong

ganeshie8 (ganeshie8):

It turns out that we cannot find \(f(x,y)\) because there is no such function whose partials with respect to \(x\) and \(y\) are \(3x^2+3y\) and \(x+2y\) respectively

OpenStudy (loser66):

Yes, I see it. Thanks for pointing it out.

OpenStudy (loser66):

How to solve my original problem and down the answer to one of the options?? Please

ganeshie8 (ganeshie8):

My point is that question should make it clear. question seems to be asking the exact value but your initial attempt focused on finding the approximation.

ganeshie8 (ganeshie8):

you've got f(0,0) = -4 done with the original prblem right ?

OpenStudy (loser66):

YOu said that it is invalid. I am just lucky. I don't want it. I want the firm logic.

ganeshie8 (ganeshie8):

Never said it is invalid.

OpenStudy (loser66):

above : "You just got lucky with the present problem because there exists a potential function, this is not the case with most functions and you will end up using approximation formula"

ganeshie8 (ganeshie8):

\(f(x,y) = x^3+3xy+y^2-4\) does satisfy all the given conditions, so we're good!

ganeshie8 (ganeshie8):

what i meant was, you cannot solve the function always. recall the definition of exact differential equations or conservative vector fields

OpenStudy (loser66):

Ok, got you. Thank you for being patient to me. :)

ganeshie8 (ganeshie8):

It seems we can find the potential function if the mixed partials are equal : \[f_{xy} = f_{yx}\] Eventhough most functions that we encounter directly satisfy this, you can see it is easy to cook up the partials that violate above condition and thus lack solution(s).